# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ Acknowledgements.
- Wang et al. (2024b) Ziliang Wang, Ge Li, Jia Li, Yihong Dong, Yingfei Xiong, and Zhi Jin. 2024b. Line-level Semantic Structure Learning for Code Vulnerability Detection. arXiv:2407.18877
- Wang et al. (2024c) Ziliang Wang, Ge Li, Jia Li, Yingfei Xiong, Jia Li, Meng Yan, and Zhi Jin. 2024c. M2CVD: Enhancing Vulnerability Semantic through Multi-Model Collaboration for Code Vulnerability Detection. arXiv:2406.05940
@@
- Xu et al. (2025) Hanxiang Xu, Shenao Wang, Ningke Li, Kailong Wang, Yanjie Zhao, Kai Chen, Ting Yu, Yang Liu, and Haoyu Wang. 2025. Large Language Models for Cyber Security: A Systematic Literature Review. arXiv:2405.04760
@@
- Yang et al. (2025a) Xu Yang, Shaowei Wang, Jiayuan Zhou, and Wenhan Zhu. 2025a. One-for-All Does Not Work! Enhancing Vulnerability Detection by Mixture-of-Experts (MoE). arXiv:2501.16454
- Yang et al. (2025b) Yixin Yang, Bowen Xu, Xiang Gao, and Hailong Sun. 2025b. Context-Enhanced Vulnerability Detection Based on Large Language Model. arXiv:2504.16877
@@
- Yildiz et al. (2025) Alperen Yildiz, Sin G. Teo, Yiling Lou, Yebo Feng, Chong Wang, and Dinil M. Divakaran. 2025. Benchmarking LLMs and LLM-based Agents in Practical Vulnerability Detection for Code Repositories. arXiv:2503.03586
@@
- Zhang et al. (2024d) Boyu Zhang, Triet H. M. Le, and M. Ali Babar. 2024d. MVD: A Multi-Lingual Software Vulnerability Detection Framework. arXiv:2412.06166
- Zhang et al. (2025a) Jie Zhang, Haoyu Bu, Hui Wen, Yongji Liu, Haiqiang Fei, Rongrong Xi, Lun Li, Yun Yang, Hongsong Zhu, and Dan Meng. 2025a. When LLMs Meet Cybersecurity: A Systematic Literature Review. Cybersecurity 8, 1 (2025), 55.
@@
- Zhang et al. (2024a) Quanjun Zhang, Chunrong Fang, Yang Xie, YuXiang Ma, Weisong Sun, Yun Yang, and Zhenyu Chen. 2024a. A Systematic Literature Review on Large Language Models for Automated Program Repair. arXiv:2405.01466
- Zhang et al. (2024b) Quanjun Zhang, Chunrong Fang, Yang Xie, Yaxin Zhang, Yun Yang, Weisong Sun, Shengcheng Yu, and Zhenyu Chen. 2024b. A Survey on Large Language Models for Software Engineering. arXiv:2312.15223
@@
- Zhang et al. (2024c) Yingying Zhang, Zhengdan Jiang, Xingkai Cheng, Han Wu, Zonghang Tian, and Shuting Lou. 2024c. A Method of SQL Injection Attack Detection Based on Large Language Models. In International Conference on Computer Network Technology and Electronic and Information Engineering (CNTEIE). IEEE, Changchun, China, 154–158. doi:10.1109/CNTEIE66268.2024.00035

### v2

@@ Acknowledgements.
- Wang et al. (2025f) Yifan Wang, Penghao Liu, and Yang Zhang. 2025f. VuL-MCBERT: A Vulnerability Detection Method Based on Self-Supervised Contrastive Learning. In International Conference on Artificial Intelligence, Big Data and Algorithms (CAIBDA). IEEE, Beijing, China, 923–927. doi:10.1109/CAIBDA65784.2025.11183103
@@
- Wang et al. (2025c) Ziliang Wang, Ge Li, Jia Li, Yihong Dong, Yingfei Xiong, and Zhi Jin. 2025c. Line-level Semantic Structure Learning for Code Vulnerability Detection. In International Conference on Internetware (Internetware). ACM, Trondheim, Norway, 269–280. doi:10.1145/3755881.3755894
- Wang et al. (2025d) Ziliang Wang, Ge Li, Jia Li, Jia Li, Meng Yan, Yingfei Xiong, and Zhi Jin. 2025d. M2CVD: Enhancing Vulnerability Understanding through Multi-Model Collaboration for Code Vulnerability Detection. ACM Transactions on Software Engineering and Methodology (TOSEM) (2025). doi:10.1145/3771923 Just Accepted.
- Wang et al. (2025e) Ziliang Wang, Ge Li, Jia Li, Hao Zhu, and Zhi Jin. 2025e. VulAgent: Hypothesis-Validation based Multi-Agent Vulnerability Detection. arXiv:2509.11523
@@
- Wei et al. (2025) Zhiyuan Wei, Xiaoxuan Yang, Jing Sun, and Zijian Zhang. 2025. Distilling Lightweight Language Models for C/C++ Vulnerabilities. arXiv:2510.06645
@@
- Wu and Zhang (2024) Jie Wu and Cheng Zhang. 2024. A Software Vulnerability Detection Model Combined with Graph Simplification. In Asia-Pacific Artificial Intelligence and Big Data Forum (AIBDF). ACM, Ganzhou Shaanxi, China, 202–207. doi:10.1145/3718491.3718525
- Wu and Xiao (2025) Yifan Wu and Lei Xiao. 2025. Sparse-MoE: Syntax-Aware Multi-view Mixture of Experts for Long-Sequence Software Vulnerability Detection. In International Conference on Advanced Data Mining and Applications (ADMA). Springer, Kyoto, Japan, 348–362. doi:10.1007/978-981-95-3456-2_24
@@
- Xu et al. (2025) Hanxiang Xu, Shenao Wang, Ningke Li, Kailong Wang, Yanjie Zhao, Kai Chen, Ting Yu, Yang Liu, and Haoyu Wang. 2025. Large Language Models for Cyber Security: A Systematic Literature Review. ACM Transactions on Software Engineering and Methodology (2025). doi:10.1145/3769676 Just Accepted.
@@
- Yang et al. (2025a) Xu Yang, Shaowei Wang, Jiayuan Zhou, and Wenhan Zhu. 2025a. One-for-All Does Not Work! Enhancing Vulnerability Detection by Mixture-of-Experts (MoE). Proceedings of the ACM on Software Engineering (PACMSE) 2, FSE, Article FSE021 (2025), 19 pages. doi:10.1145/3715736
- Yang et al. (2025b) Yixin Yang, Bowen Xu, Xiang Gao, and Hailong Sun. 2025b. Context-Enhanced Vulnerability Detection Based on Large Language Models. ACM Transactions on Software Engineering and Methodology (TOSEM) (2025). doi:10.1145/3779222 Just Accepted.
@@
- Yildiz et al. (2025) Alperen Yildiz, Sin G Teo, Yiling Lou, Yebo Feng, Chong Wang, and Dinil Mon Divakaran. 2025. Benchmarking LLMs and LLM-based Agents in Practical Vulnerability Detection for Code Repositories. In Annual Meeting of the Association for Computational Linguistics (ACL). Association for Computational Linguistics, Vienna, Austria, 30848–30865. doi:10.18653/v1/2025.acl-long.1490
@@
- Yu and Zhang (2025) Yalong Yu and Cheng Zhang. 2025. A Vulnerability Detection Framework Based on Graph Decomposition Fusion and Augmented Abstract Syntax Tree. In International Conference on Big Data, Information and Computer Network (BDICN). ACM, Guangzhou, China, 734–738. doi:10.1145/3727353.3727471
@@
- Zhang et al. (2024c) Boyu Zhang, Triet H. M. Le, and M. Ali Babar. 2024c. MVD: A Multi-Lingual Software Vulnerability Detection Framework. arXiv:2412.06166
@@
- Zhang et al. (2024a) Quanjun Zhang, Chunrong Fang, Yang Xie, Yaxin Zhang, Yun Yang, Weisong Sun, Shengcheng Yu, and Zhenyu Chen. 2024a. A Survey on Large Language Models for Software Engineering. arXiv:2312.15223
- Zhang et al. (2025a) Shaobo Zhang, Qianzhi Wang, Qin Liu, Entao Luo, and Tao Peng. 2025a. VulTrLM: LLM-assisted Vulnerability Detection via AST Decomposition and Comment Enhancement. Empirical Software Engineering (EMSE) 31, 1 (2025), 1–36.
@@
- Zhang et al. (2024b) Yingying Zhang, Zhengdan Jiang, Xingkai Cheng, Han Wu, Zonghang Tian, and Shuting Lou. 2024b. A Method of SQL Injection Attack Detection Based on Large Language Models. In International Conference on Computer Network Technology and Electronic and Information Engineering (CNTEIE). IEEE, Changchun, China, 154–158. doi:10.1109/CNTEIE66268.2024.00035
@@
- Zhu et al. (2025a) Hao Zhu, Jia Li, Cuiyun Gao, Jiaru Qian, Yihong Dong, Huanyu Liu, Lecheng Wang, Ziliang Wang, Xiaolong Hu, and Ge Li. 2025a. Specification-Guided Vulnerability Detection with Large Language Models. arXiv:2511.04014
@@
- Zhu et al. (2025b) Qiang Zhu, Zhengying Wang, Yufei Zhuang, Guo Chen, He Luo, Meiqi Wu, Hongxiang Chen, and Zhi Liu. 2025b. An Automated Code Review Framework Based on BERT and Qianwen Large Model. In IEEE International Conference on Computer Communication and Artificial Intelligence (CCAI). IEEE, Haikou, China, 960–964. doi:10.1109/CCAI65422.2025.11189422